### v1

# FrameOracle: Learning What to See and How Much to See in Videos

Abstract
Vision-language models (VLMs) have advanced video understanding, but their performance is limited by the number of input frames they can process. Existing frame sampling strategies, such as uniform or fixed-budget selection, often fail to adapt to variations in information density or task complexity, resulting in inefficiency and information loss. To address this, we present FrameOracle, a lightweight and plug-and-play module that predicts both (1) which frames are most relevant to a given query and (2) how many frames are needed. FrameOracle is trained using a four-stage curriculum, with the first three stages relying on weak proxy signals such as cross-modal similarity. In the final stage, it leverages stronger supervision from a new dataset we introduce, FrameOracle-41K, the first large-scale VideoQA collection to provide keyframe annotations specifying the minimal set of frames required to answer each question. Extensive experiments across five VLMs and six benchmarks demonstrate that FrameOracle reduces 16-frame inputs to an average of 10.4 frames without any loss in accuracy. When starting from 64-frame candidates, it reduces the input to an average of 13.9 frames while improving accuracy by 1.4%, achieving state-of-the-art efficiency-accuracy trade-offs for scalable video understanding.
1 Introduction
Rapid advances in large language models (LLMs) (Stiennon et al., 2020; Gao et al., 2022; Yang et al., 2024) have enabled vision-language models (VLMs) to integrate visual understanding with strong linguistic reasoning (Zhang et al., 2024; Bai et al., 2025; Zhang et al., 2025a). This makes VLMs highly effective for complex video tasks such as question answering (Zhang et al., 2023; Lin et al., 2024a; b; Zhao et al., 2024; Xiao et al., 2025), summarization (Hua et al., 2025; Lee et al., 2025), and instruction following (Ren et al., 2024; Qian et al., 2024). A key challenge, however, is the large volume of data these models must process. Processing every video frame is computationally expensive, making efficient frame sampling essential (Hu et al., 2025). Most VLMs currently rely on simple approaches, such as uniform sampling at a fixed frame rate or selecting a fixed number of frames. While easy to implement, these methods have clear drawbacks: in long videos, they may miss crucial information, whereas in short videos, they often introduce redundant frames that waste resources, distract the model, and obscure key moments.
To mitigate this, a growing body of work has explored keyframe selection methods (Liu et al., 2025; Park et al., 2024; Tang et al., 2025; Zhang et al., 2025d). These approaches aim to identify a subset of frames that preserves semantic content while reducing redundancy. However, most existing methods assume a fixed, preset number of keyframes, ignoring the fact that the optimal number of frames varies across videos and queries. For example, short action-centric questions (e.g., whether a ball crosses a line in sports footage) may be resolved with just a handful of frames, while long-form narrative reasoning (e.g., inferring character intentions in a film) often requires a substantially larger set of frames. A few recent methods enable adaptive frame selection, but their adaptivity remains limited. In some cases, the selector is trained jointly with the backbone VLM (Buch et al., 2025), making it non-transferable to other models. In others, adaptivity is achieved via threshold-based filtering at inference, retaining only keyframes above a preset reward threshold. While this produces variable frame counts, it is not explicitly optimized during training, reducing effectiveness and generalizability. This raises a fundamental question: How can we design a selector that identifies the most relevant frames for a given query and determines how many are needed, while generalizing across different VLMs?
To this end, we propose FrameOracle, a lightweight, plug-and-play frame selector that can be integrated with arbitrary VLMs. Unlike prior approaches that fix the number of frames in advance or require co-training with a specific backbone, FrameOracle jointly predicts (1) the importance of each frame relative to the query and (2) the number of frames to retain. The module is trained with a four-stage curriculum. The first three stages rely on weak proxy signals, such as cross-modal similarity and leave-one-out loss degradation. The final stage leverages stronger supervision from a new dataset we create, FrameOracle-41K, a large-scale VideoQA dataset with 40,992 examples and the first to provide keyframe annotations specifying the minimal frames required to answer each question. Unlike tasks such as object detection (Lin et al., 2014) or captioning (Xiong et al., 2024), no existing dataset provides ground-truth annotations identifying the keyframes. FrameOracle dynamically adapts its selections based on both video content and the prompt, operating seamlessly as a pre-processing module for any downstream VLM.
In summary, our contributions are as follows:
  • We propose FrameOracle, a lightweight and plug-and-play frame selector that dynamically predicts both which frames are most relevant and how many are needed.
  • To facilitate training, we introduce FrameOracle-41K, the first large-scale VideoQA dataset with keyframe annotations, specifying the minimal set of frames needed to answer each question.
  • We conduct extensive experiments across five VLMs and six benchmarks, showing that FrameOracle reduces 16-frame inputs to an average of 10.4 frames without any loss in accuracy. When starting from 64-frame candidates, it reduces the input to an average of 13.9 frames while improving accuracy by 1.4%, achieving state-of-the-art efficiency-accuracy trade-offs for scalable video understanding.
2 Related Work
Keyframe Selection for Video Understanding. Most existing keyframe selection methods assume a fixed frame budget: they rank candidate frames by visual–linguistic relevance or temporal salience and then retain the top-$k$ subset (Liang et al., 2024; Tan et al., 2024; Yu et al., 2025; Liu et al., 2025; Fang et al., 2025; Tang et al., 2025). Beyond this fixed-budget paradigm, some work has explored adaptive frame selection. These approaches fall into two categories. The first are agent-based methods, where large multimodal models act as decision-makers that iteratively analyze videos. For instance, VCA (Yang et al., 2025) combines curiosity-driven exploration with tree search to identify informative segments, while AKeyS (Fan et al., 2025) leverages a language agent to heuristically expand video segments and decide both which frames to retain and when to stop. However, such methods are computationally expensive due to repeated agent calls. The second category comprises approaches that require co-training with a specific VLM backbone (Buch et al., 2025; Yu et al., 2023; Guo et al., 2025), which restricts their portability. In contrast, FrameOracle is adaptive, lightweight, and model-agnostic: it learns to jointly predict which frames are relevant and how many to retain, while remaining plug-and-play across diverse VLMs.
Datasets and Supervision for Video-Language Models. Progress in video-language reasoning has been driven by large-scale datasets such as LLaVA-Video-178K (Zhang et al., 2024), ShareGPT4Video (Chen et al., 2024b), VideoRefer (Yuan et al., 2025), and CinePile (Rawal et al., 2024), which cover diverse scenarios and support both short- and long-form understanding. However, most of these datasets provide supervision only at the answer level, leaving the underlying evidence unannotated. In the absence of frame-level labels, keyframe selection methods are typically forced to rely on proxy signals, such as leave-one-out degradation or heuristic scoring. A few benchmarks, such as TVQA+ (Lei et al., 2020), ReXTime (Chen et al., 2024a), and HourVideo (Chandrasegaran et al., 2024), move toward span-level annotations, but none supply labels for both the indices of keyframes and the minimal sufficient number of frames needed to answer a question. FrameOracle-41K is the first dataset to provide explicit keyframe annotations for video–question pairs, offering high-quality supervision for both training and evaluation of adaptive frame selectors.
Figure 1: FrameOracle-41K data generation pipeline. Stage I (agent-based keyframe mining) iteratively explores each video using a multimodal agent, ultimately returning a predicted answer with confidence and relevance scores for all visited frames. Stage II (filtering and verification) first discards frames with low relevance scores and then verifies sufficiency by requiring three independent VLMs to answer correctly using only the remaining keyframes.
3 FrameOracle-41K Dataset
We introduce FrameOracle-41K, the first VideoQA dataset that provides keyframe annotations, specifying the minimal set of frames needed to answer each question. The corpus contains 40,992 video–question pairs spanning diverse scenes and durations. In contrast to existing VideoQA datasets, which provide only ground-truth answers and, in some cases, coarse temporal spans in the video, FrameOracle-41K records, for each instance, the minimal number of frames needed to answer the question with high confidence, along with the keyframes that constitute the necessary evidence. Below, we describe our data collection and generation pipeline, the verification and filtering procedures used to retain high-quality data, and the key statistics of the resulting dataset.
3.1 Data Gathering and Processing
All video–question pairs in FrameOracle-41K are sourced from LLaVA-Video-178K (Zhang et al., 2024), a large-scale VideoQA dataset that covers a wide range of scenarios and activities. From this corpus, we first select nearly 100k videos, each 2–3 minutes long, balancing adequate temporal context with a manageable annotation effort. We then apply a two-stage process to create the final dataset. Stage I (agent-based keyframe mining) iteratively explores each video using a multimodal agent. At the end of exploration, the agent outputs a predicted answer with its confidence and assigns a relevance score to every visited frame. Stage II (filtering and verification) first removes frames with relevance scores below a threshold, then verifies sufficiency by requiring three independent VLMs to answer correctly using only the remaining keyframes. Only instances that pass this verification are retained, finalizing the minimal sufficient frame count. An overview of the pipeline is shown in Figure 1. Appendix B provides an example of a data instance in JSON format.
Stage I: Agent-based Keyframe Mining. Starting from a uniformly sampled set of 64 frames, we employ an agent built on Qwen2.5-VL-72B API (Bai et al., 2025) to iteratively explore the video with respect to the given question. In the first iteration, the agent inspects three anchor frames (indices 0, 31, and 63), assigns relevance scores, and attempts an answer with a confidence estimate. It then compares the pairwise summed relevance of adjacent anchors (0+31 vs. 31+63) to decide which segment to explore next. Within the selected segment, a denser set of four anchor frames is sampled, another answer with confidence is attempted, and the same pairwise relevance comparison guides subsequent iterations. This iterative score–refine cycle continues until either the agent becomes confident enough to provide a stable answer or all frames have been examined. By the end of Stage I, the agent returns (1) its predicted answer and confidence, and (2) the complete set of frames it has inspected, each annotated with a relevance score. Any video-question pair for which the agent’s predicted answer does not match the ground-truth answer is then discarded. This mined trajectory captures both the localization of question-specific evidence and fine-grained frame-level importance signals, forming the raw candidates for the next stage.
Stage II: Filtering and Verification. After obtaining candidate keyframes from Stage I, we first remove all frames with relevance scores below a threshold $\lambda$, leaving only those with stronger relevance. For each video–question pair, we then test whether the selected keyframes alone are sufficient to answer the question. Specifically, the keyframe set and the question are fed into three independent VLMs (i.e., Qwen2.5-VL-72B (Bai et al., 2025), LLaVA-OneVision-72B (Li et al., 2025), and LLaVA-Video-72B (Zhang et al., 2024)), and their predictions are compared against the ground-truth answer. Only instances for which all three models succeed using only the keyframes are retained; otherwise, the instance is discarded. This cross-model verification ensures that the released dataset contains consistent, question-grounded keyframe annotations.
(a) Distribution of video durations in FrameOracle-41K.
(b) Distribution of the number of selected keyframes per video–question pair.
Figure 2: FrameOracle-41K dataset statistics. (a) Video durations range from 2 to 3 minutes, offering enough temporal context for analysis without including unnecessary frames. (b) Keyframe annotations are sparse: most videos require fewer than 10 frames, yet the dataset also includes harder edge cases to capture more challenging scenarios.
3.2 Dataset Statistics
Through the two-stage pipeline described above, we obtain 40,992 validated video–question pairs, each annotated with keyframes and minimal sufficient frame counts. This collection forms the FrameOracle-41K dataset. Figure 2(a) shows that most videos are 2 to 3 minutes long, long enough to provide meaningful temporal context without the inefficiency and redundancy of very long videos. Figure 2(b) shows the distribution of keyframes per video–question pair. The median is about five frames, the mean is around seven, and more than 80% of samples require no more than 10 frames. A small fraction of more challenging cases, however, require 30 or more frames. These statistics highlight the challenge and value of FrameOracle-41K: models must learn to identify minimal yet sufficient evidence across diverse temporal contexts, rather than relying on dense frame sampling.
4 Method
We introduce FrameOracle, a lightweight and adaptive frame selector that dynamically determines the appropriate number of keyframes from a video, conditioned on the user prompt. FrameOracle enables efficient video understanding by providing the downstream VLM with a compact yet highly relevant subset of frames.
Since directly processing all frames of a video, $V$, is computationally expensive, we first apply uniform temporal sampling to extract a candidate set of $N$ frames, denoted as $V_{C}=\{f_{1},\dots,f_{N}\}$. This pre-sampling step acts as a coarse filter, reducing the input to a manageable size for FrameOracle (e.g., $N=64$ or $N=16$). Our goal is to learn FrameOracle, a selection policy, $\Pi_{\theta}$, parameterized by $\theta$, that operates on the candidate set. Given a candidate set $V_{C}$ and a text prompt $P$, FrameOracle selects a compact subset of frames $V_{S}\subset V_{C}$. Unlike approaches that fix the number of selected frames in advance, FrameOracle dynamically determines the subset size, $K=|V_{S}|$, as part of the selection process. Formally, $\Pi_{\theta}$ maps the pair $(V_{C},P)$ to the selected subset $V_{S}$, which is then passed to a downstream VLM, $\mathcal{M}$, to perform a reasoning task, producing an output $A=\mathcal{M}(V_{S},P)$. The objective is to train $\Pi_{\theta}$ to choose subsets that maximize the performance of $\mathcal{M}$ while keeping $K$ as small as possible.
Figure 3: Overview of the FrameOracle pipeline. FrameOracle (dashed box) receives raw video frames and the textual prompt, and jointly predicts frame importance and the number of frames to keep. It outputs a compact keyframe subset, which is fed into the downstream VLM. $V_{C}$ denotes the pre-sampled frame collection, and $V_{S}$ denotes the subset selected by FrameOracle.
4.1 FrameOracle
FrameOracle, $\Pi_{\theta}$, is a neural module that learns to jointly predict frame importance and the number of frames to select from the candidate set $V_{C}$. We begin by extracting features from both the video frames and the text prompt, which serve as inputs to FrameOracle. For the candidate frame set $V_{C}$, we use a visual encoder to generate a sequence of $N$ frame embeddings. The text prompt $P$ is encoded using the native tokenizer of the downstream VLM. The FrameOracle architecture, shown in Figure 3, is composed of two main components: (1) a cross-modal fusion encoder and (2) dual prediction heads.
(1) Cross-Modal Fusion. To model the relationship between the text query and the video, we fuse the two modalities. Frame and text embeddings are first projected into a shared latent space using linear layers, and then processed by a stack of Transformer encoder layers. This architecture captures the complex interactions between the textual prompt and the visual content of the candidate frames.
(2) Dual Prediction Heads. The output of the fusion encoder is passed to two specialized heads, which form the core of our selection policy:
  • Rank Head: This head evaluates the relevance of each candidate frame to the prompt. It processes the fused feature sequence to output a scalar importance score, $s_{i}$, for each frame $f_{i}\in V_{C}$, resulting in a score vector $S=\{s_{1},\dots,s_{N}\}$.
  • K Head: This head predicts how many frames to select from the candidate set. It takes the globally aggregated features from the fusion encoder and outputs a probability distribution over a discrete set of possible values for $K$, where $K\leq N$.
4.2 Training
We train FrameOracle using a curriculum-based, four-stage protocol. This strategy progressively refines the policy $\Pi_{\theta}$, teaching it to reason effectively over the pre-sampled frames (e.g., 16 or 64). The staged training leverages four widely used public VideoQA datasets, covering clips ranging from roughly 10 seconds to 15 minutes in length. Details of the full dataset composition are provided in Appendix A.
Stage 1: Text-Visual Alignment. The initial stage focuses on learning a robust cross-modal representation by aligning the textual prompt with the visual content of the candidate frames. We use the pre-trained text-visual model SigLIP (Zhai et al., 2023) as a teacher to provide supervision. For each prompt-frame pair, a SigLIP similarity score serves as the target relevance signal. The feature projectors and the cross-modal Transformer encoder are trained with a RankNet loss (Burges et al., 2005), encouraging the model’s predicted scores to match the relative ordering of the SigLIP similarities. Concretely, for any two frames $i$ and $j$, let $s_{i}$ and $s_{j}$ denote their SigLIP similarity scores, and $y_{i}$ and $y_{j}$ their predicted scores. We define the pairwise preference label as $t_{ij}=\mathrm{sign}(s_{i}-s_{j})$. The RankNet loss is then given by
$\mathcal{L}_{\text{RankNet}}=\sum_{i<j}\log\!\Big(1+\exp\big(-t_{ij}\,(y_{i}-y_{j})\big)\Big),$ (1)
where $t_{ij}=0$ corresponds to tied frames and does not contribute to the gradient. In this way, the alignment capability of SigLIP is distilled into FrameOracle. The K Head remains frozen during this stage.
Stage 2: Rank Head Optimization. In the second stage, we train the Rank Head to identify the most salient frames in the candidate set $V_{C}$. Unlike the first stage, where SigLIP-based supervision is computed independently for each frame and provides no temporal guidance, this stage uses the downstream VLM’s loss as a supervisory signal, allowing the selector to capture temporal dependencies across frames. To generate training targets, we adopt a leave-one-out (LOO) approach: for each frame $f_{i}\in V_{C}$, we remove it from the input set and pass the remaining frames through the VLM, measuring the change in its loss. A larger increase indicates that $f_{i}$ is more important. These importance scores serve as soft targets, and the Rank Head is trained with a RankNet loss to predict them. During this stage, the K Head remains frozen, while the Transformer encoder and feature projectors are fine-tuned with a smaller learning rate to stabilize training.
Stage 3: K Head Optimization. The third stage focuses on training the K Head to predict the number of frames. During this stage, the Rank Head is frozen, while the feature projectors and the Transformer encoder are fine-tuned with a very small learning rate for slight adaptation. For each sample, we evaluate the downstream VLM (the same backbone and task loss as in Stage 2) using the top-$k$ frames from $V_{C}$, ranked by the frozen Rank Head, for a candidate value $k\in N$. We then select the target
$k^{*}\;=\;\arg\min_{k\in N}\Big(\mathrm{zscore}(\mathcal{L}_{\text{task}}(k))+\lambda_{k}\,k\Big),$ (2)
where the linear penalty balances accuracy and frame cost. The K Head predicts a categorical distribution $p_{\theta}(k)$ over $k\in\{1,\dots,N\}$ and is trained with
$\mathcal{L}_{K}\;=\;(1-\alpha)\,\mathcal{L}_{\text{evo}}\;+\;\alpha\,\mathcal{L}_{\text{class}},$ (3)
where the Expected Value Objective
$\mathcal{L}_{\text{evo}}\;=\;\mathrm{SmoothL1}\!\left(\sum_{k=1}^{N}k\,p_{\theta}(k),\;k^{*}\right)$
regresses the predicted expectation to $k^{*}$, and $\mathcal{L}_{\text{class}}$ is a KL divergence aligning $p_{\theta}$ with a Gaussian-shaped soft target centered at $k^{*}$.
Stage 4: Supervised Fine-tuning with Ground Truth. In the final stage, we perform supervised fine-tuning (SFT) on FrameOracle-41K, which provides supervision for both the keyframe indices and the number of frames. Unlike the earlier stages, which rely on weak or proxy signals, FrameOracle-41K offers high-quality annotations that have been verified for consistency. The Rank Head is trained to align its predictions with the annotated keyframes, while the K Head is jointly trained to match the annotated $K$ values. This strong, direct supervision further refines the selection policy beyond what is achieved in Stages 1–3.
5 Experiments
5.1 Experiment Settings
[…]
Benchmarks. We evaluate FrameOracle on six widely adopted video benchmarks, which can be divided into long-video and short-video understanding tasks. For long-video understanding, we include EgoSchema (Mangalam et al., 2023), LongVideoBench (Wu et al., 2024), MLVU (Zhou et al., 2025), and Video-MME (Fu et al., 2025), all of which require reasoning over extended temporal contexts ranging from minutes to hours. These datasets emphasize challenges such as cross-event reasoning, global consistency, and temporal grounding across lengthy sequences. For short-video understanding, we evaluate on NExTQA (Xiao et al., 2021), and Perception (Pătrăucean et al., 2023), which involve clips typically within tens of seconds. These benchmarks focus on fine-grained event recognition, local temporal relations, and reasoning within concise videos. We follow the LMMs-Eval library (Zhang et al., 2025b) for evaluation, and report accuracy across all benchmarks.
Table 1: FrameOracle vs. SOTA VLMs. “Frames” shows $M\rightarrow\bar{K}$: FrameOracle starts from $M$ uniformly sampled frames and reduces to an average of $\bar{K}$. Highlighted rows show the upper-bound performance of FrameOracle with larger frame inputs. LVB = LongVideoBench validation set.
Model Frames NExTQA Perception LVB Video-MME EgoSchema MLVU Avg.
OE_val OE_test MC
[…]
ShareGPT4Video-8B (Chen et al., 2024b) 16 - - - - 41.8 39.9 - 46.4 -
LLaMA-VID-7B (Li et al., 2024b) 16 - - - 44.6 - 25.9 38.5 33.2 -
VideoChat2-7B (Li et al., 2024a) 16 - - - - 39.3 39.5 63.6 44.5 -
VideoLLaMA2-7B (Cheng et al., 2024) 16 - - 45.4 54.9 53.1 47.9 53.1 - -
InternVL2-40B (OpenGVLab, 2024) 16 - - - - 59.3 61.2 - 59.5 -
[…]
Qwen2.5-VL-3B (Bai et al., 2025) 32 25.1 29.6 75.4 65.9 54.1 58.4 53.4 59.4 52.7
[…]
[5pt/3pt] LLaVA-OneVision-7B (Li et al., 2025) 16 14.6 16.7 78.2 56.4 55.0 56.1 60.8 60.9 49.8
[…]
[5pt/3pt] LLaVA-Video-7B (Zhang et al., 2024) 16 27.3 32.4 81.0 64.3 55.8 59.8 54.2 61.7 54.6
[…]
[5pt/3pt] VideoLLaMA3-7B (Zhang et al., 2025a) 16 27.8 32.3 82.3 72.3 56.1 61.2 61.4 50.9 55.5
[…]
[5pt/3pt] GPT-4o (Hurst et al., 2024) 16 - - 63.1 - 51.6 58.5 66.0 38.7 55.6
[…]
5.2 Results and Analysis
Comparisons with State-of-the-Art Models. Table 5.1 presents a comprehensive comparison of FrameOracle across two categories: (1) five state-of-the-art (SOTA) VLMs, and (2) its integration with five diverse VLMs, Qwen2.5-VL (Bai et al., 2025), LLaVA-OneVision (Li et al., 2025), LLaVA-Video (Zhang et al., 2024), VideoLLaMA3 (Zhang et al., 2025a), and the proprietary GPT-4o (Hurst et al., 2024). Qwen2.5-VL internally merges every two adjacent frames into a single representation. To ensure a fair comparison with models that process raw frames directly, we report the baseline using 32 frames. For each model integrated with FrameOracle, we report results for two configurations: using 16-frame FrameOracle and 64-frame FrameOracle.
Under the 16-frame condition, FrameOracle maintains accuracy comparable to the baseline models across all benchmarks while reducing the number of frames by approximately 35%. With 64-frame inputs, FrameOracle begins with a denser candidate set and adaptively selects relevant frames. In this setting, it consistently improves performance over the baseline models while still reducing frames by about 15%. This demonstrates that a larger candidate pool enables FrameOracle to better exploit temporal redundancy, resulting in improved accuracy–efficiency trade-offs. FrameOracle is trained independently and applied in a fully plug-and-play manner, requiring no co-training or backbone-specific adaptation. These results confirm its ability to generalize across model architectures.
RQ 1: Does giving a VLM more frames consistently improve its performance?
One might expect that providing more frames always improves performance, since additional frames offer more visual evidence. However, Table 5.1 shows the opposite: using more frames often fails to help and can even reduce accuracy. This aligns with recent findings that long-video reasoning is inherently sparse, with only a small subset of frames being truly relevant (Park et al., 2024). Extra frames primarily introduce redundancy and noise, diluting cross-modal attention and yielding diminishing returns (Li et al., 2023).
By contrast, when FrameOracle selects a smaller but more informative subset of frames, performance can improve, especially on open-ended benchmarks. For example, on LLaVA-OneVision-7B, reducing 16 frames to roughly 10.4 improves NExTQA metrics (OE_val: $14.6\rightarrow 16.1$, OE_test: $16.7\rightarrow 17.8$) and EgoSchema ($60.8\rightarrow 62.4$). Similar trends are observed for GPT-4o, where accuracy rises from 55.6 to 56.2 despite using fewer frames. Qualitative examples in Appendix E (Figure 6) further illustrate this effect: FrameOracle identifies the key evidence with fewer frames, yielding correct answers where naive higher-frame sampling fails.
Table 2: FrameOracle vs. SOTA keyframe selection methods. NExTQA reports MCQ. Methods using more frames or larger LLMs are shown in gray. LVB = LongVideoBench validation set.
Model Frames NExTQA LVB Video-MME EgoSchema MLVU
[…]
SeViLA (Yu et al., 2023) 8 63.6 - - 25.7 -
LVNet (Park et al., 2024) 12 72.9 - - - -
VideoAgent (Wang et al., 2024) 8.4 71.3 - - 60.2 -
FFS (Buch et al., 2025) 8.6 66.7 - - - -
MoReVQA (Min et al., 2025) 30 69.2 - - - -
VSLS (Guo et al., 2025) 32 - 63.4 63.0 - -
AKS (Tang et al., 2025) 64 - 62.7 65.3 - -
[…]
LLaVA-OneVision-7B (Li et al., 2025) 8 77.4 54.3 53.8 62.0 58.4
[…]
[5pt/3pt] LLaVA-Video-7B (Zhang et al., 2024) 8 75.6 54.2 55.9 51.8 60.5
[…]
RQ 2: How does FrameOracle compare with existing SOTA methods for keyframe selection?
We compare FrameOracle with (1) keyframe selection methods that are jointly trained with their VLM backbone in their original configurations, and (2) plug-and-play keyframe selection methods applied to open-source models (Table 5.2). The first category cannot be applied directly to open-source models, and FFS (Buch et al., 2025) is the only method that adaptively determines the number of retained frames; all other methods assume a fixed number of keyframes. All plug-and-play selection methods only provided 8-frame results. To ensure a fair comparison, we disable FrameOracle’s K Head and rely solely on the Rank Head: given 64 uniformly sampled frames, we select the top-8 ranked frames as input to the backbone VLMs.
FrameOracle achieves competitive performance compared to prior plug-and-play methods. Across NExTQA, LongVideoBench, Video-MME, and EgoSchema, it improves accuracy by roughly 2–4 percentage points, outperforming Frame-Voyager (Yu et al., 2025), BOLT (Liu et al., 2025), and KFC (Fang et al., 2025). On MLVU, FrameOracle outperforms the base VLMs but does not surpass heuristic methods such as KFC, a greedy selection strategy that maximizes relevance and diversity, which achieve higher scores. This gap reflects MLVU’s focus on fine-grained temporal grounding and multi-event reasoning, where heuristics can sometimes capture domain-specific cues more effectively. Overall, the results demonstrate that even without the K Head, the Rank Head alone can reliably prioritize important frames and deliver consistent gains across multiple VLMs, achieving state-of-the-art performance on most benchmarks.
RQ 3: How much can FrameOracle reduce computational cost while preserving accuracy?
We take LLaVA-Video-7B with 16 input frames as the baseline and report per-GPU, per-sample averages. FrameOracle reduces the input from 16 to 10.4 frames, cutting the VLM cost from 184.38 to 109.11 TFLOPs and the end-to-end total from 184.38 to 110.98 TFLOPs ($39.8\%$). It also lowers latency from 0.615 to 0.363 seconds ($41.0\%$) and reduces tokens from 11,644.0 to 7,581.6, while maintaining accuracy. With a larger candidate pool, FrameOracle reduces 64 frames to 13.9, improving accuracy by +1.4+1.4 while still lowering total compute to 167.67 TFLOPs ($9.1\%$), tokens to 10,133.1 ($13.0\%$), and latency to 0.556 seconds ($9.6\%$). These results reveal a clear trade-off: smaller pools yield larger efficiency gains with no loss in accuracy, while larger pools deliver accuracy gains with moderate compute savings. Although reported on LLaVA-Video-7B, computation is dominated by the backbone VLM, so similar efficiency–accuracy trade-offs are expected for other $\sim$7B-scale models, with only minor variation due to architectural details.
RQ 4: Are all training-stage components essential for FrameOracle’s performance?
We conduct ablations over the four training stages to evaluate the contribution of each stage, using Qwen2.5-VL-3B as the backbone VLM (Table 4). The baseline (first row) randomly selects 16 frames from 32 uniformly sampled candidates. Stage 1 (text–visual alignment) underperforms the baseline, though it provides a foundational starting point. Stages 2 (Rank Head) and 3 (K Head) yield clear performance improvements, and the full model with Stage 4 (fine-tuning on FrameOracle-41K) delivers further gains over Stages 2 and 3 on most benchmarks, with only a slight decline on Video-MME. Moreover, the average number of retained frames decreases from 21.8 to 20.9, showing that FrameOracle-41K supervision stabilizes performance while enabling higher accuracy with fewer frames. These results demonstrate that ground-truth supervision from FrameOracle-41K is essential for refining both frame importance scoring and the prediction of the number of frames, establishing it as a valuable resource for adaptive frame selection.
Table 3: Comparison of FLOPs, latency, visual tokens, and accuracy. The values of the computational cost are reported as per-GPU, per-sample averages.
Model Frames TFLOPs ↓ Latency (s) ↓ Visual Tokens ↓ Avg. Acc. ↑
DINOv2 FrameOracle VLM Total
LLaVA-Video-7B 16 184.38 184.38 0.615 11,644.0 54.6
[…]
Table 4: Four-stage training of FrameOracle, evaluated on Qwen2.5-VL-3B. Stages are added progressively to assess their impact. The baseline (first row) randomly selects 16 of 32 frames. Bold numbers indicate best performance. LVB = LongVideoBench validation set.
Model Frames NExTQA Perception LVB Video-MME EgoSchema MLVU
OE_val OE_test MC
[…]
6 Conclusion
We propose FrameOracle, a lightweight, plug-and-play frame selector that adaptively determines which frames to retain and how many are needed. To facilitate training, we introduce FrameOracle-41K, a large-scale VideoQA dataset with 40,992 examples, and the first to provide keyframe annotations specifying the minimal frames required to answer each question. Extensive experiments show that FrameOracle improves the performance of diverse VLM backbones without co-training, reducing FLOPs, latency, and the number of tokens, while also outperforming state-of-the-art keyframe selection methods. Future work will explore relaxing the fixed-length input setting to support variable-sized frame sets.
Reproducibility Statement
To support reproducibility, we provide details on both the model and dataset. FrameOracle’s design, including its learning objective, selection policy, and staged curriculum, is described in Section 4, with training procedures and hyperparameters in Appendix A (Figure 4; Table 5). FrameOracle-41K construction, including agent-based mining, verification, and dataset format, is covered in Section 3 and Appendix B. Evaluation settings, including benchmarks, backbones, and metrics, are in Section 5.1. These sections provide all the information needed to reproduce our results.
Ethics Statement
FrameOracle-41K is created from source videos collected from the internet, which may contain images of individuals and reflect societal biases present in online content. Our data processing pipeline does not involve identifying or profiling any individuals. The data is used solely for developing our video understanding model. We release the dataset strictly for non-commercial, academic research purposes and caution future users to be aware of potential inherent biases in the data.
Acknowledgments
This research was partially supported by the National Eye Institute (NEI) of the National Institutes of Health (NIH) under award number R01EY034562. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix
Appendix A Full Implementation Details
Figure 4: Four-stage training strategy of FrameOracle. The model is progressively optimized from weak to strong supervision, culminating in supervised fine-tuning with FrameOracle-41K annotations. Fire icons indicate trainable modules, while snowflake icons denote frozen ones.
Table 5: Datasets used in FrameOracle training. Stage 4 leverages FrameOracle-41K.
Task Dataset Amount
Stage1 LLaVA-Video-178K (Zhang et al., 2024), ShareGPT4o-Video (Chen et al., 2024b), Video-ChatGPT (Maaz et al., 2024) 300K
Stage2 LLaVA-Video-178K (Zhang et al., 2024), LLaVA-Hound (Zhang et al., 2025c), Video-ChatGPT (Maaz et al., 2024) 300K
Stage3 LLaVA-Video-178K (Zhang et al., 2024), LLaVA-Hound (Zhang et al., 2025c), Video-ChatGPT (Maaz et al., 2024) 300K
Stage4 FrameOracle-41K (Our Dataset) 40K
Training strategy illustration. Figure 4 presents a schematic of our four-stage curriculum, highlighting trainable modules (fire) and frozen modules (snowflake) at each stage.
[…]
Datasets used in staged training. FrameOracle is optimized using a four-stage curriculum with progressively stronger supervision. Stages 1-3 rely on large-scale video–language corpora, while Stage 4 leverages our FrameOracle-41K dataset. Table 5 summarizes the dataset composition for each stage.
Stage 1: Cross-modal alignment. K Head is frozen while the feature projectors and cross-modal Transformer encoder are trained jointly, both optimized with a learning rate of $1\times 10^{-4}$. The 16-frame selector uses a batch size of 16 and trains for approximately 48 hours, whereas the 64-frame version uses a smaller batch size of 2 and completes in about 91 hours.
Stage 2: Rank Head optimization. Rank Head is trained while the K Head remains frozen. The Rank Head uses a learning rate of $1\times 10^{-4}$, and the feature projectors and Transformer encoder are fine-tuned with a smaller learning rate of $1\times 10^{-5}$. The 16-frame selector uses a batch size of 16 and trains for approximately 40 hours, whereas the 64-frame variant uses the same batch size and takes about 52 hours.
Stage 3: K Head optimization. K Head is the primary trainable module, optimized with a learning rate of $1\times 10^{-4}$. The feature projectors and Transformer encoder are lightly updated with a learning rate of $1\times 10^{-7}$, while the Rank Head remains frozen. We set $\lambda_{k}=0.0105$ to balance accuracy and efficiency. The 16-frame selector uses a batch size of 16 and trains for approximately 35 hours, whereas the 64-frame variant uses the same batch size and takes about 60 hours.
Stage 4: Supervised fine-tuning on FrameOracle-41K. Rank Head and K Head are trained jointly with a learning rate of $5\times 10^{-5}$, while the feature projectors and Transformer encoder are fine-tuned with $1\times 10^{-5}$. The 16-frame selector is trained with a batch size of 8 for approximately 12 hours, and the 64-frame version uses the same batch size and trains for about 18 hours.
Appendix B FrameOracle-41K Data Format
We release the FrameOracle-41K dataset in JSON format, with each entry corresponding to a single video–question pair. Each entry includes the instance identifier, question–answer pair, paths to the associated video and extracted keyframes, video duration, and number of selected frames. Below, we provide an example JSON entry to illustrate the dataset’s structure.
"id": 30,
"question": "What folding technique is demonstrated first in the video?",
"ground_truth_answer": "The ’SHIKAKU NO GI’ (Square Fold) technique is demonstrated first.",
"video": "/srv/nfs/video_data/video/ytb_8yhoV5C3bT8.mp4",
"keyframes_dir": "/srv/nfs/video_data/extracted_frames/ytb_8yhoV5C3bT8",
"duration": 126.893,
"num_selected_frames": 8
Appendix C Prompts for Data Generation
Prompt Template for Stage I: Initial Frame Analysis You are analyzing a video that is {duration_seconds} seconds long. The video has been uniformly sampled into 64 frames, indexed from 0 (start) to 63 (end). Analyze these {len(initial_indices)} initial frames (indices: {initial_indices}) to answer: “{question}”. Provide a short caption for each frame, a relevance score (INTEGER 1-5), your confidence (high/medium/low), and your answer attempt. Respond in JSON: {{“frame_analysis”: [{{“index”: int, “caption”: “str”, “relevance”: int}}], “confidence”: “str”, “answer_attempt”: “str”, “reasoning”: “str”}} IMPORTANT GUIDELINES: - Relevance combines BOTH (a) how well the frame’s TEMPORAL POSITION matches the question mentioned, and (b) how much the visible CONTENT answers the question. A high score (4-5) requires strong evidence on both axes. - You may use “high” confidence early ONLY IF: You have seen explicit, definitive evidence that unquestionably answers the question (e.g., clearly visible target object/person/action). - Before setting “high” confidence, explicitly mention in your reasoning: (a) Why current evidence is sufficient. (b) Why additional unseen frames are unlikely to alter your conclusion. - If there’s any reasonable scenario where unseen frames could alter your answer, you must explicitly acknowledge that and keep your confidence at “medium”. Follow these instructions strictly.
Prompt Template for Stage II: Deep-dive Analysis and Refinement You are analyzing a video that is {duration_seconds} seconds long. The video has been uniformly sampled into 64 frames, indexed from 0 (start) to 63 (end). Current context on question “{question}”: Current context in buffer “{buffer}”. Now analyze these {len(indices)} new frames (indices: {[int(idx) for idx in indices]}) from the gap ({start_idx}, {end_idx}). Tasks: - Provide a caption, relevance score (INTEGER 1-5) for each NEW frame, your UPDATED confidence, answer, and reasoning. - If the new evidence changes your view of any PREVIOUS frame listed above, list the updated scores under “revised_prev_scores” (index, new relevance 1-5). Respond in JSON: {{“new_frame_analysis”: [{{“index”: int, “caption”: “str”, “relevance”: int}}], “revised_prev_scores”: [{{“index”: int, “relevance”: int}}], “confidence”: “str”, “answer_attempt”: “str”, “reasoning”: “str”}} IMPORTANT GUIDELINES: - Relevance combines BOTH (a) how well the frame’s TEMPORAL POSITION matches the question mentioned, and (b) how much the visible CONTENT answers the question. A high score (4-5) requires strong evidence on both axes. - You may use “high” confidence early ONLY IF: You have seen explicit, definitive evidence that unquestionably answers the question (e.g., clearly visible target object/person/action). - Before setting “high” confidence, explicitly mention in your reasoning: (a) Why current evidence is sufficient. (b) Why additional unseen frames are unlikely to alter your conclusion. - If there’s any reasonable scenario where unseen frames could alter your answer, you must explicitly acknowledge that and keep your confidence at “medium”. Follow these instructions strictly.
Appendix D Additional Benchmarks Details
Table 6 summarizes the evaluation prompts for each benchmark used in our experiments, most of which are adapted from LMMs-Eval.
Table 6: Prompts specifying the response format used for each evaluation benchmark.
Benchmark Response formatting prompts
MLVU
Video-MME Answer with the option’s letter from the given choices directly.
EgoSchema Answer with the option’s letter from the given choices directly.
NExTQA
Perception Answer with the option’s letter from the given choices directly.
LongVideoBench Answer with the option’s letter from the given choices directly.
Appendix E Qualitative Examples
As shown in Figure 5, FrameOracle can achieve the same correct answers as uniform sampling while using far fewer frames. In the illustrated examples, our selector retained only 4 frames or, in some cases, just 2 frames out of the original 16 inputs, yet still provided sufficient evidence to answer the questions correctly. This demonstrates that many uniformly sampled frames are redundant and that FrameOracle effectively filters them out without compromising accuracy.
Figure 6 shows cases aligned with RQ1 (see Section 5.2). Feeding all 16 uniformly sampled frames can mislead the VLM with irrelevant or distracting content, producing incorrect answers. In contrast, FrameOracle selects a smaller, query-aligned subset, enabling the model to focus on relevant evidence and provide accurate answers. These examples demonstrate that more frames do not guarantee better performance, and adaptive selection of fewer, informative frames improves understanding.
Figure 5: Qualitative examples. FrameOracle answers correctly while using only a few frames (2 to 4 out of 16), compared to uniform sampling, which relies on the full input.
Figure 6: Qualitative examples for RQ1. Using all 16 uniformly sampled frames can produce incorrect answers, whereas FrameOracle answers correctly by selecting only the relevant subset.